The decimal value of the number stored in 01000000010000000000000000000000 is

3

-2.5

10.5

10.01
Answer #1

Answer : 3

Explanation :

According to IEEE 754 floating point number to decimal conversion is like below:
The given floating point number is 0100 0000 0100 0000 0000 0000 0000 0000
Sign = 1 bit (0)
Exponent = 8 bits (10000000)
Mantissa = 23 bits (100000000000000000000000)

Formula: (1)^sign 2^exponent-127 (1*Mantissa)
sign = 1^1 = "+"
Exponent = 10000000 = 1*2^7 = 128
Mantissa = 1*2^1 = 0.5

Applying formula:
=2^(128-127) [1 + 0.5]

= 2^1 1.5

= 3

So the Answer is : 3

Please upvote if you find this answer helpful, thanks in advance!